A 20kg traveler's bag is being pulled along the floor with a force of 50N. The force is applied on the handle that forms an angle of 30
1
Expert's answer
2021-09-27T10:12:00-0400

Is the force is applied at 30° above the horizontal and we neglect friction, the acceleration will be


"a=\frac{F_\text{net}}{m}=\frac{F\cos\theta}{m}=2.2\text{ m/s}^2."